# ContextCache: Context-Aware Semantic Cache for Multi-Turn Queries in Large Language Models

### v1

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at XXXXXX.
@@ 1. Introduction
To address these challenges, we present ContextCache, a context-aware semantic caching framework for multi-turn conversational systems. Our approach overcomes existing limitations through three key innovations. First, our Dynamic Context Modeling implements a hierarchical self-attention mechanism that captures cross-turn semantic dependencies, mitigating attention dilution in concatenation while preserving turn-specific features lost through representation flattening. Second, our LLM-enhanced training employs difficult negative sample mining to improve matching precision, addressing insufficient semantic discrimination through robust sentence-level supervision. Finally, our Two-Stage Dynamic Retrieval Architecture combines efficient vector retrieval for candidate selection with precise attention-based contextual matching, balancing computational efficiency with semantic accuracy. Experimental evaluation shows ContextCache outperforms GPTCache, improving precision and recall by 12.9% and 15.7% while cache-served responses deliver approximately 10 times lower latency compared to LLM invocation. Our contributions include:
@@ 3.3. Core Technique And Evaluation
Context integration. Our self-attention mechanism models semantic relationships across dialogue turns, generating unified contextual representations that capture inter-turn dependencies. Figure 3(a) demonstrates that our approach significantly outperforms concatenation and averaging methods by preserving turn-specific semantic features while modeling their contextual relationships. Compared to GPTCache, this context-aware approach achieves 12.9% higher precision and 15.7% improved recall, particularly in distinguishing between semantically similar queries that appear in different conversational contexts.

### v2

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 14, No. 12 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/uYanJX/ContextCache.
@@ 1. Introduction
To address these challenges, we present ContextCache, a context-aware semantic caching framework for multi-turn conversational systems. Our approach overcomes existing limitations through three key innovations. First, our Dynamic Context Modeling implements a hierarchical self-attention mechanism that captures cross-turn semantic dependencies, mitigating attention dilution in concatenation while preserving turn-specific features lost through representation flattening. Second, our LLM-enhanced training employs difficult negative sample mining to improve matching precision, addressing insufficient semantic discrimination through robust sentence-level supervision. Finally, our Two-Stage Dynamic Retrieval Architecture combines efficient vector retrieval for candidate selection with precise attention-based contextual matching, balancing computational efficiency with semantic accuracy. Experimental evaluation shows ContextCache outperforms GPTCache, improving precision and recall by 10.9% and 14.8% while cache-served responses deliver approximately 10 times lower latency compared to LLM invocation. Our contributions include:
@@ 3.3. Core Technique And Evaluation
Context integration. Our self-attention mechanism models semantic relationships across dialogue turns, generating unified contextual representations that capture inter-turn dependencies. Figure 3(a) demonstrates that our approach significantly outperforms concatenation and averaging methods by preserving turn-specific semantic features while modeling their contextual relationships. Compared to GPTCache, this context-aware approach achieves 10.9% higher precision and 14.8% improved recall, particularly in distinguishing between semantically similar queries that appear in different conversational contexts.